Solve the problem.
A store sells tents,sleeping bags, and camp stools. A customer buys a tent, 5 sleeping bags, and 4 camp stools for $242. The price of the tent is 9 times the cost of a camp stool. The cost of a sleeping bag is $16 more than the cost of a camp stool. Find the cost of each item.
I have tried using this x=tents, y=sleeping bags, and z=camp stools.
x+5y+4z=242 and then I am stuck, please help.

Answer by scott8148(6628)

good so far...

"The price of the tent is 9 times the cost of a camp stool" ___ x = 9z

"The cost of a sleeping bag is $16 more than the cost of a camp stool" ___ y = z + 16

substituting ___ 9z + 5(z + 16) + 4z = 242 ___ 9z + 5z + 80 + 4z = 242

18z = 162 ___ z = 9

substituting ___ y = 9 + 16 = 25

substituting ___ x = 9(9) = 81